# Converting Transformers into DGNNs Form

### v1

@@ 1 Introduction
Therefore, replacing self-attention has become another well-known option (Tay et al., 2021a; Lee-Thorp et al., 2022; Yu et al., 2022a). However, achieving high-rank or full-rank attention matrices while maintaining a time complexity lower than quadratic is a challenge. We address this challenge via synthetic digraph convolution  [^1] In this work, directed graphs are abbreviated as digraphs.. Since self-attention is closely related to digraph convolution (Zaheer et al., 2020), why not replace self-attention with digraph convolution? This work is based on this hypothesis. We synthesize a unitary digraph convolution called Synvolution, which achieves high performance while maintaining linear time complexity for long sequences. We also apply the kernel polynomial method (Silver and Röder, 1994; Wang, 1994; Wang and Zunger, 1994; Vijay et al., 2004; Weiße et al., 2006; Weiße and Fehske, 2008) as an alternative technique for the multi-head operation. We refer to Synvolution with the kernel polynomial method as Kernelution. Supported by the theoretical foundation of the kernel polynomial method, the filter of Kernelution can function as any corresponding unitary filter required by distinct datasets.
@@
  • We propose Synvolution, a novel self-attention alternative with linear time complexity.
@@ 3.1 Self-Attention and Multi-Head Self-Attention
where $\tau=\sqrt{D_{h}}$ is the temperature parameter, the softmax function is applied row-wise, $\mathrm{T}$ represents the transpose operation, $\mathbf{X}\mathbf{W}_{\text{Q}}\in\mathbb{R}^{N\times D_{h}}$, $\mathbf{X}\mathbf{W}_{\text{K}}\in\mathbb{R}^{N\times D_{h}}$, and $\mathbf{X}\mathbf{W}_{\text{V}}\in\mathbb{R}^{N\times D_{h}}$ are referred to as the query, key, and value matrix, in which $\mathbf{W}_{\text{Q}}\in\mathbb{R}^{D\times D_{h}}$, $\mathbf{W}_{\text{K}}\in\mathbb{R}^{D\times D_{h}}$, and $\mathbf{W}_{\text{V}}\in\mathbb{R}^{D\times D_{h}}$ are learnable parameters.
Conventionally, the Multi-Head Self-Attention (MHSA) with $H$ heads is commonly chosen to enhance performance. It is defined as
@@ 4.1 Synvolution
$\centering\begin{bmatrix}G_{ii}&G_{ji}\\ G_{ij}&G_{jj}\end{bmatrix}=\begin{bmatrix}\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu(\frac{\alpha+\beta}{2})}\cos{(\frac{\gamma}{2})}&-\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu(\frac{\alpha-\beta}{2})}\sin{(\frac{\gamma}{2})}\\ \mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu(\frac{\alpha-\beta}{2})}\sin{(\frac{\gamma}{2})}&\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu(\frac{\alpha+\beta}{2})}\cos{(\frac{\gamma}{2})}\end{bmatrix},\@add@centering$ (6)
@@
Here, $\mathbf{H}^{(l)}_{\text{l}}\in\mathbb{C}^{N\times N}$ is a lower unitary Hessenberg matrix, $\mathbf{H}^{(l)}_{\text{u}}\in\mathbb{C}^{N\times N}$ is an upper unitary Hessenberg matrix, and $\mathbf{P}^{(l)}\in\mathbb{R}^{N\times N}$ is a permutation matrix that either fixed (Dao et al., 2022), learnable (Mena et al., 2018), or even an identity matrix $\mathbf{I}_{N}\in\mathbb{R}^{N\times N}$.
@@
Given an input signal $\mathbf{x}\in\mathbb{C}^{N}$ and an output signal $\mathbf{y}\in\mathbb{C}^{N}$, the time complexity of $L$-DHHP as a discrete unitary transform with the fast implementation as $\mathbf{y}=\bm{\Phi}\mathbf{x}$ is $\mathcal{O}(LN)$.
@@
[…]
@@ 4.2.2 Kernel Polynomial Method
This results in an intuitive penalty applied to the Chebyshev coefficients, with higher order Chebyshev coefficients incurring greater penalties than the lower ones. It causes the Chebyshev polynomial interpolation with the kernel polynomial loss to simulate the kernel polynomial method with a dynamic, learnable kernel. We apply the kernel polynomial method with Synolution, which turns out what we call Kernelution. The corresponding formula is defined as
$\centering\mathrm{Kern}(\mathbf{X})=\bm{\Phi}^{-1}\left[\exp\left(\mathrm{i}\mkern 1.0mu\cdot{p}_{\text{KP}}(\mathbf{\Lambda})\right)\odot(\bm{\Phi}\mathbf{X}\mathbf{W}_{\text{V}})\right].\@add@centering$ (12)
@@ 4.3 Gated Feed-Forward Network and PostScaleNorm
$\centering\mathrm{GFFN}(\mathbf{X})=\left[\mathrm{sp}(\Re(\mathbf{X})\mathbf{W}_{\Re})\odot\tanh(\Im(\mathbf{X})\mathbf{W}_{\Im})\right]\mathbf{W}_{\text{O}},\@add@centering$ (13)
where $\mathrm{sp}(\cdot)$ represents the $\mathrm{softplus}$ function (Dugas et al., 2000), $\mathbf{W}_{\Re}\in\mathbb{R}^{D\times D_{\text{hid}}}$, $\mathbf{W}_{\Im}\in\mathbb{R}^{D\times D_{\text{hid}}}$ and $\mathbf{W}_{\text{O}}\in\mathbb{R}^{D_{\text{hid}}\times D}$ are trainable weight matrices. We let the real part to learn the magnitude, and the imaginary part to learn the sign. Besides, we apply the PostNorm architecture (Wang et al., 2019) with ScaleNorm (Nguyen and Salazar, 2019) across the whole model, namely PostScaleNorm. Specifically, we apply $\zeta\cdot\Re(\mathbf{Z})+(1-\zeta)\cdot\Im(\mathbf{Z})$ for complex-valued signal $\mathbf{Z}$, where $\zeta\in[0,1]$ is a learnable parameter.
@@ 5.2 Long Document Classification
Model LongDoc16K LongDoc32K
@@ 5.3 DNA Sequence-based Taxonomy Classification
Model Ensembl (B/S) Ensembl (M/R)
@@ 5.4 Ablation Studies
Method ListOps Text Retrieval Image Pathfinder
@@ 6 Conclusion and Future Work
In this work, we introduce Converter, a Transformer variant that replaces self-attention with a synthetic unitary digraph convolution called Synvolution. By leveraging the inverse process of eigendecomposition and LQ factorization, we synthesize a unitary digraph shift operator with learnable eigenvalues and eigenvectors. Our fast $1$-DHHP implementation achieves linear time complexity while preserving the full-rank property of Synvolution. We further enhance Synvolution by incorporating the kernel polynomial method, resulting in Kernelution. We propose a kernel polynomial loss to enable dynamic kernel adaptation during training.
@@ Appendix A Pseudocode for Synvolution and Kernelution
def dhhp_trans(transform=True, x, gl_ii, gl_ij, gl_ji, gl_jj, gu_ii, gu_ij, gu_ji, gu_jj, diag):
y = torch.zeros_like(x) # b, n, d
z = torch.zeros_like(x) # b, n, d
if (transform is True) and (m is not None):
@@
elif (transform is False) and (diag is not None):
@@
# First row (j = 0)
y[:,0,:] = gu_ii[:,0].unsqueeze(-1) * x[:,0,:] + gu_ij[:,0].unsqueeze(-1) * x[:,1,:]
# Middle rows (j in [1, n-2])
y[:,1:n-1,:] = (gu_ji[:,:-1].unsqueeze(-1) * x[:,:-2,:] + (gu_jj[:,:-1] * gu_ii[:,1:]).unsqueeze(-1) * x[:,1:n-1,:] + (gu_jj[:,:-1] * gu_ij[:,1:]).unsqueeze(-1) * x[:,2:,:])
# Last row (j = n-1)
y[:,n-1,:] = gu_ji[:,n-2].unsqueeze(-1) * x[:,n-2,:] + gu_jj[:,n-2].unsqueeze(-1) * x[:,n-1,:]
# First row (i = 0)
z[:,0,:] = gl_ii[:,0] * y[:,0,:] + gl_ij[:,0] * y[:,1,:]
# Middle rows (i in [1, n-2])
z[:,1:n-1,:] = ((gl_ii[:,1:] * gl_ji[:,:-1]).unsqueeze(-1) * y[:,:-2,:] + (gl_ii[:,1:] * gl_jj[:,:-1]).unsqueeze(-1) * y[:,1:n-1,:] + gl_ij[:,1:].unsqueeze(-1) * y[:,2:,:])
# Last row (i = n-1)
z[:,n-1,:] = gl_ji[:,n-2].unsqueeze(-1) * y[:,n-2,:] + gl_jj[:,n-2].unsqueeze(-1) * y[:,n-1,:]
if (transform is True) and (diag is not None):
@@
elif (transform is False) and (m is not None):
z = z.reshape(b, n // m, m, d).transpose(1, 2).reshape(b, n, d)
@@
def inverse_dhhp_trans(transform=False, x, gu_ii_conj_trs, gu_ij_conj_trs, gu_ji_conj_trs, gu_jj_conj_trs, gl_ii_conj_trs, gl_ij_conj_trs, gl_ji_conj_trs, gl_jj_conj_trs, diag_conj_trs):
return dhhp_trans(transform, x, gu_ii_conj_trs, gu_ij_conj_trs, gu_ji_conj_trs, gu_jj_conj_trs, gl_ii_conj_trs, gl_ij_conj_trs, gl_ji_conj_trs, gl_jj_conj_trs, diag_conj_trs)
@@
g_ij = torch.exp(-1j * (alpha - beta) / 2) * torch.sin(gamma / 2)
g_ji = -torch.exp(1j * (alpha - beta) / 2) * torch.sin(gamma / 2)
@@
def kernelution(eigenvalue, K, g, mu, x, alpha, beta, gamma, theta):
@@
eigenvalue = kernel_polynomial_method(eigenvalue, K, mu, g)
g_ii, g_ij, g_ji, g_jj = givens_rot_para(alpha, beta, gamma)
diag = torch.exp(1j * theta)
gu_ii_conj_trs, gu_ij_conj_trs, gu_ji_conj_trs, gu_jj_conj_trs, gl_ii_conj_trs, gl_ij_conj_trs, gl_ji_conj_trs, gl_jj_conj_trs = givens_rot_para_conj_trs(g_ii, g_ij, g_ji, g_jj)
@@
x = dhhp_trans(True, x, gl_ii, gl_ij, gl_ji, gl_jj, gu_ii, gu_ij, gu_ji, gu_jj, diag)
@@
x = inverse_dhhp_trans(False, x, gu_ii_conj_trs, gu_ij_conj_trs, gu_ji_conj_trs, gu_jj_conj_trs, gl_ii_conj_trs, gl_ij_conj_trs, gl_ji_conj_trs, gl_jj_conj_trs, diag_conj_trs)

### v3

@@ 1 Introduction
Therefore, replacing self-attention has become another well-known option (Tay et al., 2021a; Lee-Thorp et al., 2022; Yu et al., 2022a). However, achieving high-rank or full-rank attention matrices while maintaining a time complexity lower than quadratic is a challenge. We address this challenge via synthetic digraph convolution  [^1] In this work, directed graphs are abbreviated as digraphs.. Since self-attention is closely related to digraph convolution (Zaheer et al., 2020), why not replace self-attention with digraph convolution? This work is based on this hypothesis. We synthesize a unitary digraph convolution called Synvolution, which achieves high performance while maintaining linearithmic time complexity for long sequences. We also apply the kernel polynomial method (Silver and Röder, 1994; Wang, 1994; Wang and Zunger, 1994; Vijay et al., 2004; Weiße et al., 2006; Weiße and Fehske, 2008) as an alternative technique for the multi-head operation. We refer to Synvolution with the kernel polynomial method as Kernelution. Supported by the theoretical foundation of the kernel polynomial method, the filter of Kernelution can function as any corresponding unitary filter required by distinct datasets.
@@
  • We propose Synvolution, a novel self-attention alternative with linearithmic time complexity.
@@ 3.1 Self-Attention and Multi-Head Self-Attention
where $\tau=\sqrt{D_{h}}$ is the temperature parameter, the softmax function is applied row-wise, $\mathrm{T}$ represents the transpose operation, $\mathbf{X}\mathbf{W}_{\text{Q}}\in\mathbb{R}^{N\times D_{h}}$, $\mathbf{X}\mathbf{W}_{\text{K}}\in\mathbb{R}^{N\times D_{h}}$, and $\mathbf{X}\mathbf{W}_{\text{V}}\in\mathbb{R}^{N\times D_{h}}$ are referred to as the query, key, and value matrix, in which $\mathbf{W}_{\text{Q}}\in\mathbb{R}^{D\times D_{h}}$, $\mathbf{W}_{\text{K}}\in\mathbb{R}^{D\times D_{h}}$, and $\mathbf{W}_{\text{V}}\in\mathbb{R}^{D\times D_{h}}$ are learnable parameters. Conventionally, the Multi-Head Self-Attention (MHSA) with $H$ heads is commonly chosen to enhance performance. It is defined as
@@ 4.1 Synvolution
$\centering\begin{bmatrix}G_{ii}&G_{ij}\\ G_{ji}&G_{jj}\end{bmatrix}=\begin{bmatrix}\overline{c}&-s\\ \overline{s}&c\end{bmatrix}=\begin{bmatrix}\mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu(\frac{\alpha+\beta}{2})}\cos{(\frac{\gamma}{2})}&-\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu(\frac{\alpha-\beta}{2})}\sin{(\frac{\gamma}{2})}\\ \mathrm{e}\mkern 1.0mu^{-\mathrm{i}\mkern 1.0mu(\frac{\alpha-\beta}{2})}\sin{(\frac{\gamma}{2})}&\mathrm{e}\mkern 1.0mu^{\mathrm{i}\mkern 1.0mu(\frac{\alpha+\beta}{2})}\cos{(\frac{\gamma}{2})}\end{bmatrix},\@add@centering$ (6)
@@
Here, $\mathbf{H}^{(l)}_{\text{l}}\in\mathbb{C}^{N\times N}$ is a lower unitary Hessenberg matrix, $\mathbf{H}^{(l)}_{\text{u}}\in\mathbb{C}^{N\times N}$ is an upper unitary Hessenberg matrix, and $\mathbf{P}^{(l)}\in\mathbb{R}^{N\times N}$ is a permutation matrix that either learnable (Mena et al., 2018), fixed (Dao et al., 2022), or even an identity matrix $\mathbf{I}_{N}\in\mathbb{R}^{N\times N}$.
@@
Given an input signal $\mathbf{x}\in\mathbb{C}^{N}$ and an output signal $\mathbf{y}\in\mathbb{C}^{N}$, the time complexity of $L$-DHHP as a discrete unitary transform with the fast implementation as $\mathbf{y}=\bm{\Phi}\mathbf{x}$ is $\mathcal{O}(LN\log{N})$.
@@
$\centering\mathrm{Synv}(\mathbf{X}\mathbf{W}_{\text{V}})=\bm{\Phi}^{-1}\left[\exp(\mathrm{i}\mkern 1.0mu\mathbf{\Lambda})\odot(\bm{\Phi}\mathbf{X}\mathbf{W}_{\text{V}})\right],\@add@centering$ (8)
@@ 4.2.2 Kernel Polynomial Method
This results in an intuitive penalty applied to the Chebyshev coefficients, with higher order Chebyshev coefficients incurring greater penalties than the lower ones. It causes the Chebyshev polynomial interpolation with the kernel polynomial loss to simulate the kernel polynomial method with a learnable kernel. We apply the kernel polynomial method with Synolution, which turns out what we call Kernelution. The corresponding formula is defined as
$\centering\mathrm{Kern}(\mathbf{X}\mathbf{W}_{\text{V}})=\bm{\Phi}^{-1}\left[\exp\left(\mathrm{i}\mkern 1.0mu\cdot{p}_{\text{KP}}(\mathbf{\Lambda})\right)\odot(\bm{\Phi}\mathbf{X}\mathbf{W}_{\text{V}})\right].\@add@centering$ (12)
@@ 4.3 Gated Feed-Forward Network and PostScaleNorm
$\centering\mathrm{GFFN}(\mathbf{X})=\left[\mathrm{softplus}(\Re(\mathbf{X})\mathbf{W}_{\Re})\odot\tanh(\Im(\mathbf{X})\mathbf{W}_{\Im})\right]\mathbf{W}_{\text{O}},\@add@centering$ (13)
where $\mathbf{W}_{\Re}\in\mathbb{R}^{D\times D_{\text{hid}}}$, $\mathbf{W}_{\Im}\in\mathbb{R}^{D\times D_{\text{hid}}}$ and $\mathbf{W}_{\text{O}}\in\mathbb{R}^{D_{\text{hid}}\times D}$ are trainable weight matrices. We let the real part to learn the magnitude, and the imaginary part to learn the sign. Besides, we apply the PostNorm architecture (Wang et al., 2019) with ScaleNorm (Nguyen and Salazar, 2019) across the whole model, namely PostScaleNorm. Specifically, we apply $\mathrm{ScaleNorm}(\mathbf{Z}+\zeta\cdot\Re(\mathbf{Z})+(1-\zeta)\cdot\Im(\mathbf{Z}))$ for a complex-valued signal $\mathbf{Z}$, where $\zeta\in[0,1]$ is a learnable parameter.
@@ 5.2 Long Document Classification
Model LongDoc16K $\uparrow$ LongDoc32K $\uparrow$
@@ 5.3 DNA Sequence-based Taxonomy Classification
Model Ensembl (B/S) $\uparrow$ Ensembl (M/R) $\uparrow$
@@ 5.4 Ablation Studies
Method ListOps $\uparrow$ Text $\uparrow$ Retrieval $\uparrow$ Image $\uparrow$ Pathfinder $\uparrow$
@@ 6 Conclusion and Future Work
In this work, we introduce Converter, a Transformer variant that replaces self-attention with a synthetic unitary digraph convolution called Synvolution. By leveraging the inverse process of eigendecomposition and LQ factorization, we synthesize a unitary digraph shift operator with learnable eigenvalues and eigenvectors. Our fast $1$-DHHP implementation achieves linearithmic time complexity while preserving the full-rank property of Synvolution. We further enhance Synvolution by incorporating the kernel polynomial method, resulting in Kernelution. We propose a kernel polynomial loss to enable dynamic kernel adaptation during training.
@@ Appendix A Pseudocode for Synvolution and Kernelution
Algorithm 1 PyTorch-like pseudocode for Parallel Scan.
# b: batch size, n: length, d: feature dimension
class PScan(torch.autograd.Function):
@staticmethod
def expand_(A, X):
if A.size(1) == 1:
return
T = 2 * (A.size(1) // 2)
Aa = A[:, :T].view(A.size(0), T//2, 2, -1)
Xa = X[:, :T].view(X.size(0), T//2, 2, -1)
Xa[:, :, 1].add_(Aa[:, :, 1] * Xa[:, :, 0])
Aa[:, :, 1].mul_(Aa[:, :, 0])
PScan.expand_(Aa[:, :, 1], Xa[:, :, 1])
Xa[:, 1:, 0].add_(Aa[:, 1:, 0] * Xa[:, :-1, 1])
Aa[:, 1:, 0].mul_(Aa[:, :-1, 1])
if T < A.size(1):
X[:, -1].add_(A[:, -1] * X[:, -2])
A[:, -1].mul_(A[:, -2])
@staticmethod
def acc_rev_(A, X):
if X.size(1) == 1:
return
T = 2 * (X.size(1) // 2)
Aa = A[:, -T:].view(A.size(0), T//2, 2, -1)
Xa = X[:, -T:].view(X.size(0), T//2, 2, -1)
Xa[:, :, 0].add_(Aa[:, :, 1].conj() * Xa[:, :, 1])
B = Aa[:, :, 0].clone()
B[:, 1:].mul_(Aa[:, :-1, 1].conj())
PScan.acc_rev_(B, Xa[:, :, 0])
Xa[:, :-1, 1].add_(Aa[:, 1:, 0].conj() * Xa[:, 1:, 0])
if T < A.size(1):
X[:, 0].add_(A[:, 1].conj() * X[:, 1])
@staticmethod
def forward(ctx, A, X, Y_init):
ctx.A = A[:, :, None].clone()
ctx.Y_init = Y_init[:, None, :].clone()
ctx.A_star = ctx.A.clone()
ctx.X_star = X.clone()
PScan.expand_(ctx.A_star, ctx.X_star)
return ctx.A_star * ctx.Y_init + ctx.X_star
@staticmethod
def backward(ctx, grad_output):
U = grad_output * ctx.A_star.conj()
A = ctx.A.clone()
R = grad_output.clone()
PScan.acc_rev_(A, R)
Q = ctx.Y_init.expand_as(ctx.X_star).clone()
Q[:, 1:].mul_(ctx.A_star[:, :-1].conj()).add_(ctx.X_star[:, :-1])
grad_A = (Q.conj() * R).sum(-1)
return grad_A, R, U.sum(dim=1)
pscan = PScan.apply
@@
def dhhp_trans(transform=True, x, g_l_ii, g_l_ij, g_l_ji, g_l_jj, g_u_ii, g_u_ij, g_u_ji, g_u_jj, diag):
y = torch.zeros_like(x)
z = torch.zeros_like(x)
if transform is True:
@@
else:
@@
x_, y = torch.zeros_like(x), torch.zeros_like(x)
x_[:, :-1, :] = g_u_ii.unsqueeze(-1) * x[:, :-1, :]
x_ = x_.flip(1)
g_u_ij = g_u_ij.flip(1)
g_u_ij = torch.cat([g_u_ij, torch.zeros_like(g_u_ij[:, :1])], dim=1)
p_u = torch.zeros_like(g_u_ij)
p_u[:, 1:] = g_u_ij[:, :-1].clone()
h_u_init = x_[:, 0, :].clone()
h_u = pscan(p_u, x_, h_u_init)
h_u = h_u.flip(1)
y[:, 1:, :] = g_u_ji.unsqueeze(-1) * x[:, :-1, :] + g_u_jj.unsqueeze(-1) * h_u[:, 1:, :]
y[:, 0, :] = h_u[:, 0, :]
y_, z = torch.zeros_like(y), torch.zeros_like(y)
y_[:, 1:, :] = g_l_jj.unsqueeze(-1) * y[:, 1:, :]
g_l_ji = torch.cat([g_l_ji, torch.zeros_like(g_l_ji[:, :1])], dim=1)
p_l = torch.zeros_like(g_l_ji)
p_l[:, 1:] = g_l_ji[:, :-1].clone()
h_l_init = y_[:, 0, :].clone()
h_l = pscan(p_l, y_, h_l_init)
z[:, :-1, :] = g_l_ii.unsqueeze(-1) * h_l[:, :-1, :] + g_l_ij.unsqueeze(-1) * y[:, 1:, :]
z[:, n-1, :] = h_l[:, n-1, :]
if transform is True:
@@
else:
z = z.reshape(b, m, n // m, d).transpose(1, 2).reshape(b, n, d)
@@
def inverse_dhhp_trans(transform=False, x, g_l_ii_conj_trs, g_l_ij_conj_trs, g_l_ji_conj_trs, g_l_jj_conj_trs, g_u_ii_conj_trs, g_u_ij_conj_trs, g_u_ji_conj_trs, g_u_jj_conj_trs, diag_conj_trs):
return dhhp_trans(transform, x, g_u_ii_conj_trs, g_u_ij_conj_trs, g_u_ji_conj_trs, g_u_jj_conj_trs, g_l_ii_conj_trs, g_l_ij_conj_trs, g_l_ji_conj_trs, g_l_jj_conj_trs, diag_conj_trs)
@@
g_ij = -torch.exp(1j * (alpha - beta) / 2) * torch.sin(gamma / 2)
g_ji = torch.exp(-1j * (alpha - beta) / 2) * torch.sin(gamma / 2)
@@
def kernelution(eigenvalue, K, g, mu, x, alpha_l, beta_l, gamma_l, alpha_u, beta_u, gamma_u, theta):
@@
eigenvalue = kernel_polynomial_method(eigenvalue, K, g, mu)
g_l_ii, g_l_ij, g_l_ji, g_l_jj = givens_rot_para(alpha_l, beta_l, gamma_l)
g_u_ii, g_u_ij, g_u_ji, g_u_jj = givens_rot_para(alpha_u, beta_u, gamma_u)
diag = torch.exp(2j * math.pi * theta)
g_l_ii_conj_trs, g_l_ij_conj_trs, g_l_ji_conj_trs, g_l_jj_conj_trs = givens_rot_para_conj_trs(g_l_ii, g_l_ij, g_l_ji, g_l_jj)
g_u_ii_conj_trs, g_u_ij_conj_trs, g_u_ji_conj_trs, g_u_jj_conj_trs = givens_rot_para_conj_trs(g_u_ii, g_u_ij, g_u_ji, g_u_jj)
@@
x = dhhp_trans(True, x, g_l_ii, g_l_ij, g_l_ji, g_l_jj, g_u_ii, g_u_ij, g_u_ji, g_u_jj, diag)
@@
x = inverse_dhhp_trans(False, x, g_l_ii_conj_trs, g_l_ij_conj_trs, g_l_ji_conj_trs, g_l_jj_conj_trs, g_u_ii_conj_trs, g_u_ij_conj_trs, g_u_ji_conj_trs, g_u_jj_conj_trs, diag_conj_trs)